OpenStudy (anonymous):

Evaluate 48 ÷ (6 + 2) • 3 4 18 14 30

OpenStudy (anonymous):

18

OpenStudy (anonymous):

so your work

OpenStudy (anonymous):

show*

OpenStudy (anonymous):

It's multiple choice ^

OpenStudy (anonymous):

and what does that have to do with anything

OpenStudy (anonymous):

Evaluate x − 3y − 3z when x = −4, y = 1, and z = −1 −10 10 4 −4

OpenStudy (anonymous):

-4 - (3)(1) - (3)(-1) -4 - 3 + 3 -7 + 3 = -4

OpenStudy (anonymous):

one question at a time

OpenStudy (anonymous):

Evaluate 3x − 5y when x = 3 and y = −2 −19 19 1 −1

OpenStudy (anonymous):

-_-

OpenStudy (anonymous):

what ? lol

OpenStudy (anonymous):

nothing nothing

OpenStudy (anonymous):

lol

OpenStudy (anonymous):

3(3) - 5(-2)= 9 - (-10)= 9+10= 19

OpenStudy (anonymous):

i should get two medals now

OpenStudy (anonymous):

lol , help me with one more

OpenStudy (anonymous):

alright i will

OpenStudy (anonymous):

Rose worked the following problem incorrectly. Part 1) In which line was the mistake made? Part 2) Show your work and give the correct answer. Line 1 30 ÷ 5(3 − 1) − 4 Line 2 30 ÷ 5(2) − 4 Line 3 30 ÷ 10 − 4 Line 4 3 − 4 Line 5 −1

OpenStudy (anonymous):

is this multiple choice

OpenStudy (anonymous):

@heysa1121

OpenStudy (anonymous):

no

OpenStudy (anonymous):

you are supouse to say where did they make the mistake

OpenStudy (anonymous):

i cant help u then. i only help with multiple choice

OpenStudy (anonymous):

jk

OpenStudy (anonymous):

oh lol i was about to say

OpenStudy (anonymous):

line 3 is where the mistake is

OpenStudy (anonymous):

and how do i correct it ? lol

OpenStudy (anonymous):

patience

OpenStudy (anonymous):

30 / 5(3-1) - 4 30 / 5(2) - 4 6(2) - 4 12-4=8

OpenStudy (anonymous):

there u go

OpenStudy (anonymous):

Thank you lol i just suck at math

OpenStudy (anonymous):

yw and don't say that about yourself

OpenStudy (anonymous):

im sure youll get better

OpenStudy (anonymous):

like i dont understand it all an im doing flvs for it ans im failing

OpenStudy (anonymous):

what class r u taking

OpenStudy (anonymous):

algebra 1

OpenStudy (anonymous):

i took that on flvs too. have u tryed khanacademy it helped me out

OpenStudy (anonymous):

no i havent

OpenStudy (anonymous):

ill give u a link

OpenStudy (anonymous):

https://www.khanacademy.org/

OpenStudy (anonymous):

message me it

OpenStudy (anonymous):

to late

OpenStudy (anonymous):

thanks for fanning i knew you wouldnt be able to resist the puppy